OpenStudy (anonymous):

Type the first equation from the group you selected and identify the slope and the y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Type the second equation from the group you selected and change the equation into slope-intercept form. Be sure to show all of your work. Identify the slope and y-intercept of the equation. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. In a paragraph of at least three sentences, explain how the graphs

OpenStudy (anonymous):

Help please?

OpenStudy (reemii):

what group of equations?

OpenStudy (anonymous):

negative three-halvesx + 4 –3x + 2y = 8 y = negative 2 over 3x – 2 –2x + 3y = -6 y = negative 1 over 3x + 3 –x + 3y = 9 y = negative 1 over 6x + 5 –x + 6y = 30

OpenStudy (anonymous):

I would like to do the second one.

OpenStudy (anonymous):

If that's okay.

OpenStudy (reemii):

y = negative 2 over 3x – 2 –2x + 3y = -6 that?

OpenStudy (reemii):

and is "y = negative 2 over 3x – 2" the same as \(y=-\frac23x-2\)

OpenStudy (anonymous):

Yes! That's right!

OpenStudy (anonymous):

Yep! I am having a problem with the second equation

OpenStudy (reemii):

\(-2x+3y=-6\); the thing to know is, where to go? >> The slope and intercept can be read when the equation is in hte fomr \(y=ax+b\). Therefore, you must make your equation look like that. -> isolate the variable \(y\). do you get something?

OpenStudy (anonymous):

I'm sorry?

OpenStudy (reemii):

I meant: do you obtain* something? some results?

OpenStudy (anonymous):

Yes but I felt uncertain of what I was doing and stopped

OpenStudy (reemii):

now, what is your equation when you have isolated \(y\) ?

OpenStudy (phi):

the first step, starting with \[ -2x+3y=-6 \] is to add +2x to both sides, like this \[ +2x -2x+3y=-6 +2x\] now simplify, what do you get ?

OpenStudy (phi):

on the left side of the =, you have 2 x's take away 2 x's. How many x's do you have left ?

OpenStudy (anonymous):

One?

OpenStudy (phi):

if you have 2 dollars and take away 2 dollars, how many dollars do you have? if you have 2 cows, and take away 2 cows, how many cows do you have ? if you have 2 x's, and take away 2 x's, how many x's do you have ?

OpenStudy (anonymous):

Oh I'm sorry! none!

OpenStudy (phi):

so \[ +2x−2x+3y=−6+2x \] can be simplified to \[ 3y = 2x - 6 \] (we can switch the order of -6+2x to 2x + -6 or just 2x - 6) now divide both sides (all terms by 3) can you do that ?

OpenStudy (anonymous):

Yes! Is it...

OpenStudy (anonymous):

y = 2/3 - 2

OpenStudy (phi):

yes, but you left out the x

OpenStudy (anonymous):

Take you so much Phi!